OpenStudy (anonymous):

A cylinder is to be made so that its volume is equal to that of a sphere with a radius of 3 inches. If a cylinder is to have a height of 4 inches, find its radius.

OpenStudy (anonymous):

pi(r^2)(h)=4/3(pi)r^3 cylinder=sphere pi(r^2)(4)=pi(3)^3(4/3) pi's cancel r^2(4)=3^2(4) 4's cancel r^2=3^2 r=3

OpenStudy (jamesj):

Vol(sphere) = 4pi/3.r^3 = 4pi/3.3^3 = 36pi [r here is radius of the sphere] Vol(cylinder) = pi.r^2.h = pi.r^2.4 [r here is the radius of the base of the cylinder] If Vol(sphere) = Vol(cylinder) then 4pi.r^2 = 36pi => r^2 = 9 and hence r = 3.

OpenStudy (anonymous):

Thanks